# Characterizing Graph Datasets for Node Classification: Homophily–Heterophily Dichotomy and Beyond

### v4

@@ Acknowledgements
We thank Yao Ma for sharing the semi-synthetic datasets from Ma et al. (2022) that we used in Section 4.2. We also thank Andrey Ploskonosov for thoughtful discussions.
@@ Appendix A Random configuration model
Numerous random graph models have been proposed to reflect and predict important quantitative and topological aspects of real-world networks Boccaletti et al. [2006]. The simplest model is the Erdős–Rényi random graph, i.e., we assume that $G$ is sampled uniformly at random from the set of all graphs with $n$ nodes and $|E|$ edges. However, this model is known to be not a good descriptor of real-world networks since its Poisson degree distribution significantly differs from heavy-tailed degree distributions observed in real-world networks. A standard solution is to consider a random graph with a given degree sequence Molloy and Reed [1995]: a graph is sampled uniformly from the set of all graphs with a given degree sequence.
configuration model, which we assume throughout the paper, is defined as follows. To generate a graph, we form a set $A$ containing $d(v)$ distinct copies of each node $v$ and then choose a random matching of the elements of $A$. In this case, self-loops and multiple edges may appear. Under some conditions, the obtained graph is simple (i.e., does not contain self-loops and multiple edges) with probability $1-o(1)$ Molloy and Reed [1995].
Let us also note that there is another model that is similar to the one discussed above but can be easier to analyze. To obtain a graph with given expected degree sequence Chung and Lu [2002], we take the degree sequence from the observed graph and say that the number of edges between $i$ and $j$ follows a Poisson distribution with the mean $\frac{d(i)d(j)}{2|E|}$ if $i\neq i$ and the expected number of self-loops for a node $i$ is $\frac{d(i)^{2}}{4|E|}$. This model does not preserve the exact degree sequence but has it in expectation. Note that usually $d(i)d(j)\ll 2|E|$, so multiple edges rarely appear. Asymptotically, this model is similar to the configuration model, but a graph with a given expected degree sequence is easier to analyze theoretically. In our analysis, we assume the configuration model, so we have to track the error terms carefully.
@@ B.1 Class homophily
Recall that class homophily is defined as [Lim et al., 2021]:
@@ B.2 Constant baseline
There are several possible ways to formalize the constant baseline property. In the literature Gösgens et al. [2021a, b], this property is often formalized as follows: assuming some randomized model, the expected value of a measure should be equal to some constant. For homophily measures, this corresponds to the following definition.
@@
This property is very strict as minor variations in the definition of the model (e.g., different alternatives of the configuration model discussed in Appendix A) may lead to negligibly small error terms preventing us from getting exactly the same constant $c_{base}$. Thus, in the main text, we use the alternative Definition 2. This definition is weaker in the sense that we allow asymptotically negligible deviations from the constant $c_{base}$. On the other hand, our definition is somewhat stronger since we also require the concentration of the value around its expectation. In that sense, our definition is stronger than the asymptotic constant baseline defined in Gösgens et al. [2021a].
@@ B.3 Monotonicity property
First, let us revisit how monotonicity is defined for measures used in other areas. In general, the monotonicity property is defined as follows: if an object $B^{\prime}$ is definitely more similar to an object $A$ than $B$, then the similarity between $B^{\prime}$ and $A$ should be larger than between $B$ and $A$. The only problem is to define what it means to be “definitely more similar”. In Gösgens et al. [2021b], the monotonicity property was introduced for cluster similarity measures. Such measures evaluate how close two partitions of the same set of items are. To formally define the concept of more similar partitions, Gösgens et al. [2021b] use the concept of perfect splits and perfect merges — such transformations of a partition $B$ that make it more similar to $A$. Later, monotonicity was defined for classification evaluation measures that compare the predicted labels for a set of items with the reference ones Gösgens et al. [2021a]. Here it is easier to formalize what it means that a labeling $B^{\prime}$ is definitely more similar to the reference labeling than $B$. Indeed, if we consider a labeling $B$ and change one incorrect label to the correct one, then it becomes closer to the reference labeling.
@@
In this paper, we follow the edge rewiring approach, and thus our definition aligns well with edge-wise homophily measures and with previous works on classification and clustering performance measures Gösgens et al. [2021a, b].
If we consider edge-wise measures (that can be defined in terms of the class adjacency matrix) and directly follow the definition of monotonicity in Gösgens et al. [2021a, b], our definition would be as follows: a homophily measure is monotone if it increases when we decrement $c_{ij}$ and $c_{ji}$ by one and increment either $c_{ii}$ or $c_{jj}$ by two for $i\neq j$. This condition corresponds to taking an edge with $y_{u}\neq y_{v}$ and changing $y_{u}$ to $y_{v}$ or vice versa.
However, an important property that has to be taken into account when discussing monotonicity for homophily measures is the fact that such measures are used to compare different graph datasets, in contrast to classification evaluation measures that compare the agreement between predicted labelings with a fixed given reference labeling. This means that for classification measures, monotonicity and constant baseline are critical for fixed numbers of classes and elements. In contrast, homophily measures have to be comparable across datasets with different sizes and numbers of classes. For instance, balanced accuracy has a constant baseline with the expected value of $\frac{1}{C}$, which is sufficient for classification evaluation but is a drawback for measuring homophily. Because of this, we do not consider the standard definition of monotonicity as it is restricted to a fixed number of classes only. Instead, we introduce tolerance to empty classes, and based on that, we introduce a stronger version of monotonicity where we allow separate edge additions or deletions. In fact, our definition of monotonicity (Definition 4) is similar to strong monotonicity in Gösgens et al. [2021a, b], but also requires empty class tolerance since comparing measures across different numbers of classes is crucial.
@@ B.5 How homophily and modularity are related
Modularity is arguably the most well-known measure of goodness of a partition for a graph. It was first introduced in Newman and Girvan [2004] and is widely used in community detection literature: modularity is directly optimized by some algorithms, used as a stopping criterion in iterative methods, or used as a metric to compare different algorithms when no ground truth partition is available. The basic idea is to consider the fraction of intra-community edges among all edges of $G$ and penalize it for avoiding trivial partitions like those consisting of only one community of size $n$. In its general form and using the notation adopted in this paper, modularity is
@@ Appendix D Additional related work
In the aspect of characterizing graph datasets, our work is conceptually similar to a recent paper by Liu et al. [2022]. In this paper, the authors empirically analyze what aspects of a graph dataset (e.g., node features or graph structure) influence the GNN performance. Liu et al. [2022] follow a data-driven approach and measure the performance change caused by several data perturbations. In contrast, we follow a theoretical approach to choosing graph characteristics based only on the label-structure relation. Additionally, our characteristics are very simple, model agnostic, and can be used for general-purpose graph analysis (beyond graph ML). Similarly to Liu et al. [2022], we believe that the proposed characteristics can help in the selection and development of diverse future graph benchmarks.
@@ E.1 Datasets
Cora, citeseer, and pubmed Giles et al. [1998], McCallum et al. [2000], Sen et al. [2008], Namata et al. [2012], Yang et al. [2016] are three classic paper citation network benchmarks. For cora and citeseer labels correspond to paper topics, while for pubmed labels specify the type of diabetes addressed in the paper. Coauthor-cs and coauthor-physics Shchur et al. [2018] are co-authorship networks. Nodes represent authors, and two nodes are connected by an edge if the authors co-authored a paper. Node labels correspond to fields of study. Amazon-computers and amazon-photo Shchur et al. [2018] are co-purchasing networks. Nodes represent products, and an edge means that two products are frequently bought together. Labels correspond to product categories. Lastfm-asia is a social network of music streaming site LastFM users who live in Asia Rozemberczki and Sarkar [2020]. Edges represent follower relationships, and labels correspond to user’s nationality. In the facebook Rozemberczki et al. [2021] graph nodes correspond to official Facebook pages, and links indicate mutual likes. Labels represent site categories. In the github Rozemberczki et al. [2021] graph, nodes represent GitHub users and edges represent follower relationships. A binary label indicates that a user is either a web or a machine learning developer. Ogbn-arxiv and Ogbn-products Hu et al. [2020] are two datasets from the recently proposed Open Graph Benchmark. Ogbn-arxiv is a citation network graph with labels corresponding to subject areas. ogbn-products is a co-purchasing network with labels corresponding to product categories. Actor Tang et al. [2009], Pei et al. [2020] is a popular dataset for node classification in heterophilous graphs. The nodes correspond to actors and edges represent co-occurrence on the same Wikipedia page. The labels are based on words from an actor’s Wikipedia page. Flickr Zeng et al. [2020] is a graph of images with labels corresponding to image types. Deezer-europe Rozemberczki and Sarkar [2020] is a user network of the music streaming service Deezer with labels representing a user’s gender. Twitch-de and twitch-pt Rozemberczki et al. [2021] are social network graphs of gamers from the streaming service Twitch. [^5] Twitch-de and twitch-pt are subgraphs of a larger dataset twitch-gamers. We report characteristics for all of them since they have different sizes, edge density, and may have different structural properties. The labels indicate if a streamer uses explicit language. Genius Lim and Benson [2021], twitch-gamers Rozemberczki and Sarkar [2021], arxiv-year Hu et al. [2020], snap-patents Leskovec et al. [2005], and wiki Lim et al. [2021] are recently proposed large-scale heterophilous datasets. For the wiki dataset we remove all isolated nodes. Roman-empire, amazon-ratings, minesweeper, workers, and questions Platonov et al. [2023] are recently proposed mid-scale heterophilous datasets. We additionally construct one more binary classification graph — twitter-hate. This graph is based on data from Hateful Users on Twitter dataset on Kaggle. [^6] https://www.kaggle.com/datasets/manoelribeiro/hateful-users-on-twitter The labels indicate if a user posts hateful messages or not. We remove all the unlabeled nodes from the graph and use the largest connected component of the resulting graph.
@@ E.2 Dataset characteristics
As for label informativeness, it can be seen that on real heterophilous datasets it is typically very low (with the exception of roman-empire dataset). This is in contrast to synthetic datasets used for experiments in Luan et al. [2022], Ma et al. [2022], which sometimes exhibit a combination of low homophily and high label informativeness. High label informativeness of these datasets leads to strong GNN performance on them despite low homophily levels.
@@ F.1 Experimental setup
We use GCN Kipf and Welling [2017], GraphSAGE Hamilton et al. [2017], GAT Veličković et al. [2018] and Graph Transformer (GT) Shi et al. [2020] as representative GNN architectures for our experiments. For GraphSAGE, we use the version with the mean aggregation function and do not use the node sampling technique used in the original paper. For all models, we add a two-layer MLP after every graph neighborhood aggregation layer and further augment all models with skip connections He et al. [2016], layer normalization Ba et al. [2016], and GELU activation functions Hendrycks and Gimpel [2016]. For all models, we use two graph neighborhood aggregation layers and hidden dimension of 512. We use Adam Kingma and Ba [2015] optimizer with a learning rate of $3\cdot 10^{-5}$ and train for 1000 steps, selecting the best step based on the validation set performance. We use a dropout probability of 0.2 during training. Our models are implemented using PyTorch Paszke et al. [2019] and DGL Wang et al. [2019].
@@ F.2 Synthetic data based on SBM model
In Figure 2 of the main text, we show the results for GraphSAGE. In Figure 3, we additionally plot the dependence of GraphSAGE accuracy on homophily and label informativeness. It can be seen that the model’s accuracy is much more correlated with label informativeness than with homophily. The results for GCN, GAT, and GT are presented in Figures F.4, F.4, and F.4, respectively. As can be seen, the performance of all models generally follows the same pattern, and $\mathrm{LI}$ is a better predictor of model performance than homophily. In particular, when $\mathrm{LI}$ is high, all models achieve high accuracy even if homophily is negative. Recall that we provide Spearman correlation coefficients between model accuracy and adjusted homophily or $\mathrm{LI}$ in Table 3.
F.3 Semi-synthetic data from Ma et al. [2022]
We use the same experimental setting described above for training GCN, GraphSAGE, GAT, and GT on modifications of the cora and citeseer graphs from Ma et al. [2022]. The results for GraphSAGE and cora are provided in Figure 2 of the main text. The results for GraphSAGE and citeseer are shown in Figure F.4.
@@ F.4 Synthetic data based on LFR model
Figure 7: Accuracy of GraphSAGE on semi-synthetic citeseer graphs from Ma et al. [2022]
F.5 Synthetic data from Luan et al. [2022]
@@
Figure 8: Label informativeness depending on edge homophily on synthetic graphs from Luan et al. [2022]
The authors investigate how different levels of homophily affect GNN performance. They find that the curve showing the dependence of GNN performance on edge homophily (as well as on node homophily) is $U$-shaped: GNNs show strong results not only when edge homophily is high, but also when edge homophily is very low. Our label informativeness explains this behavior. We use the same data generating process as in Luan et al. [2022] and find that the curve of label informativeness depending on edge homophily is also $U$-shaped (see Figure 8). Thus, on the datasets from Luan et al. [2022], GNNs perform well exactly when label informativeness is high regardless of edge homophily. The $U$-shape of the label informativeness curve is not surprising: when edge homophily is very low, knowing that a node has a neighbor of a certain class provides us with information that this node probably does not belong to this class.
F.6 Correlation of GNN performance with homophily and LI on real datasets
Analyzing which measures better agree with GNN performance on real data is non-trivial. Indeed, if we consider diverse node classification problems, there are many factors affecting the GNN performance: the size of the datasets and fraction of training examples, the edge density, the number of features and their informativeness, the number of classes and their balance, and so on. However, as requested by the reviewers, we analyze how well different homophily measures and LI correlate with GNN performance on the datasets in Table 5. In this experiment, we measure the ROC AUC score since some of the datasets are highly unbalanced, and thus accuracy is not a suitable performance measure. For the datasets with more than two classes, we report the macro-averaged ROC AUC score. The relation between ROC AUC and LI for GAT and GraphSAGE is shown in Figure 9. The Spearman correlation for all the measures is reported in Table 7. We see that the largest correlation is achieved by $\mathrm{LI}_{edge}$ and $\mathrm{LI}_{node}$. Let us note that this table does not aim to compare homophily measures with each other since a good homophily measure is expected to properly capture the tendency of edges to connect similar nodes. Thus, we compare homophily measures using the proposed theoretical framework in Section 2.
(a) GAT
(b) GraphSAGE
Figure 9: Performance of GAT and GraphSAGE on real graphs vs $\mathrm{LI}_{edge}$
Table 7: Spearman correlation coefficients between ROC AUC score and characteristics of real-world datasets from Table 5
Model $h_{edge}$ $h_{node}$ $h_{class}$ $h_{adj}$ $\mathrm{LI}_{edge}$ $\mathrm{LI}_{node}$
GraphSAGE 0.64 0.62 0.67 0.61 0.78 0.82
GAT 0.69 0.65 0.74 0.69 0.85 0.84

### v5

@@ Acknowledgements
We thank Yao Ma for sharing the semi-synthetic datasets from Ma et al. [24] that we used in Section 4.2. We also thank Andrey Ploskonosov for thoughtful discussions.
@@ Appendix A Random configuration model
Numerous random graph models have been proposed to reflect and predict important quantitative and topological aspects of real-world networks [3]. The simplest model is the Erdős–Rényi random graph, i.e., we assume that $G$ is sampled uniformly at random from the set of all graphs with $n$ nodes and $|E|$ edges. However, this model is known to be not a good descriptor of real-world networks since its Poisson degree distribution significantly differs from heavy-tailed degree distributions observed in real-world networks. A standard solution is to consider a random graph with a given degree sequence [26]: a graph is sampled uniformly from the set of all graphs with a given degree sequence.
configuration model, which we assume throughout the paper, is defined as follows. To generate a graph, we form a set $A$ containing $d(v)$ distinct copies of each node $v$ and then choose a random matching of the elements of $A$. In this case, self-loops and multiple edges may appear. Under some conditions, the obtained graph is simple (i.e., does not contain self-loops and multiple edges) with probability $1-o(1)$ [26].
Let us also note that there is another model that is similar to the one discussed above but can be easier to analyze. To obtain a graph with given expected degree sequence [5], we take the degree sequence from the observed graph and say that the number of edges between $i$ and $j$ follows a Poisson distribution with the mean $\frac{d(i)d(j)}{2|E|}$ if $i\neq i$ and the expected number of self-loops for a node $i$ is $\frac{d(i)^{2}}{4|E|}$. This model does not preserve the exact degree sequence but has it in expectation. Note that usually $d(i)d(j)\ll 2|E|$, so multiple edges rarely appear. Asymptotically, this model is similar to the configuration model, but a graph with a given expected degree sequence is easier to analyze theoretically. In our analysis, we assume the configuration model, so we have to track the error terms carefully.
@@ B.1 Class homophily
Recall that class homophily is defined as [20]:
@@ B.2 Constant baseline
There are several possible ways to formalize the constant baseline property. In the literature [8, 9], this property is often formalized as follows: assuming some randomized model, the expected value of a measure should be equal to some constant. For homophily measures, this corresponds to the following definition.
@@
This property is very strict as minor variations in the definition of the model (e.g., different alternatives of the configuration model discussed in Appendix A) may lead to negligibly small error terms preventing us from getting exactly the same constant $c_{base}$. Thus, in the main text, we use the alternative Definition 2. This definition is weaker in the sense that we allow asymptotically negligible deviations from the constant $c_{base}$. On the other hand, our definition is somewhat stronger since we also require the concentration of the value around its expectation. In that sense, our definition is stronger than the asymptotic constant baseline defined in [8].
@@ B.3 Monotonicity property
First, let us revisit how monotonicity is defined for measures used in other areas. In general, the monotonicity property is defined as follows: if an object $B^{\prime}$ is definitely more similar to an object $A$ than $B$, then the similarity between $B^{\prime}$ and $A$ should be larger than between $B$ and $A$. The only problem is to define what it means to be “definitely more similar”. In [9], the monotonicity property was introduced for cluster similarity measures. Such measures evaluate how close two partitions of the same set of items are. To formally define the concept of more similar partitions, Gösgens et al. [9] use the concept of perfect splits and perfect merges — such transformations of a partition $B$ that make it more similar to $A$. Later, monotonicity was defined for classification evaluation measures that compare the predicted labels for a set of items with the reference ones [8]. Here it is easier to formalize what it means that a labeling $B^{\prime}$ is definitely more similar to the reference labeling than $B$. Indeed, if we consider a labeling $B$ and change one incorrect label to the correct one, then it becomes closer to the reference labeling.
@@
In this paper, we follow the edge rewiring approach, and thus our definition aligns well with edge-wise homophily measures and with previous works on classification and clustering performance measures [8, 9].
If we consider edge-wise measures (that can be defined in terms of the class adjacency matrix) and directly follow the definition of monotonicity in [8, 9], our definition would be as follows: a homophily measure is monotone if it increases when we decrement $c_{ij}$ and $c_{ji}$ by one and increment either $c_{ii}$ or $c_{jj}$ by two for $i\neq j$. This condition corresponds to taking an edge with $y_{u}\neq y_{v}$ and changing $y_{u}$ to $y_{v}$ or vice versa.
However, an important property that has to be taken into account when discussing monotonicity for homophily measures is the fact that such measures are used to compare different graph datasets, in contrast to classification evaluation measures that compare the agreement between predicted labelings with a fixed given reference labeling. This means that for classification measures, monotonicity and constant baseline are critical for fixed numbers of classes and elements. In contrast, homophily measures have to be comparable across datasets with different sizes and numbers of classes. For instance, balanced accuracy has a constant baseline with the expected value of $\frac{1}{C}$, which is sufficient for classification evaluation but is a drawback for measuring homophily. Because of this, we do not consider the standard definition of monotonicity as it is restricted to a fixed number of classes only. Instead, we introduce tolerance to empty classes, and based on that, we introduce a stronger version of monotonicity where we allow separate edge additions or deletions. In fact, our definition of monotonicity (Definition 4) is similar to strong monotonicity in [8, 9], but also requires empty class tolerance since comparing measures across different numbers of classes is crucial.
@@ B.5 How homophily and modularity are related
Modularity is arguably the most well-known measure of goodness of a partition for a graph. It was first introduced in [30] and is widely used in community detection literature: modularity is directly optimized by some algorithms, used as a stopping criterion in iterative methods, or used as a metric to compare different algorithms when no ground truth partition is available. The basic idea is to consider the fraction of intra-community edges among all edges of $G$ and penalize it for avoiding trivial partitions like those consisting of only one community of size $n$. In its general form and using the notation adopted in this paper, modularity is
@@ Appendix D Additional related work
In the aspect of characterizing graph datasets, our work is conceptually similar to a recent paper by Liu et al. [21]. In this paper, the authors empirically analyze what aspects of a graph dataset (e.g., node features or graph structure) influence the GNN performance. Liu et al. [21] follow a data-driven approach and measure the performance change caused by several data perturbations. In contrast, we follow a theoretical approach to choosing graph characteristics based only on the label-structure relation. Additionally, our characteristics are very simple, model agnostic, and can be used for general-purpose graph analysis (beyond graph ML). Similarly to [21], we believe that the proposed characteristics can help in the selection and development of diverse future graph benchmarks.
@@ E.1 Datasets
Cora, citeseer, and pubmed [6, 25, 38, 27, 46] are three classic paper citation network benchmarks. For cora and citeseer labels correspond to paper topics, while for pubmed labels specify the type of diabetes addressed in the paper. Coauthor-cs and coauthor-physics [39] are co-authorship networks. Nodes represent authors, and two nodes are connected by an edge if the authors co-authored a paper. Node labels correspond to fields of study. Amazon-computers and amazon-photo [39] are co-purchasing networks. Nodes represent products, and an edge means that two products are frequently bought together. Labels correspond to product categories. Lastfm-asia is a social network of music streaming site LastFM users who live in Asia [35]. Edges represent follower relationships, and labels correspond to user’s nationality. In the facebook [37] graph nodes correspond to official Facebook pages, and links indicate mutual likes. Labels represent site categories. In the github [37] graph, nodes represent GitHub users and edges represent follower relationships. A binary label indicates that a user is either a web or a machine learning developer. Ogbn-arxiv and Ogbn-products [14] are two datasets from the recently proposed Open Graph Benchmark. Ogbn-arxiv is a citation network graph with labels corresponding to subject areas. ogbn-products is a co-purchasing network with labels corresponding to product categories. Actor [42, 32] is a popular dataset for node classification in heterophilous graphs. The nodes correspond to actors and edges represent co-occurrence on the same Wikipedia page. The labels are based on words from an actor’s Wikipedia page. Flickr [47] is a graph of images with labels corresponding to image types. Deezer-europe [35] is a user network of the music streaming service Deezer with labels representing a user’s gender. Twitch-de and twitch-pt [37] are social network graphs of gamers from the streaming service Twitch. [^5] Twitch-de and twitch-pt are subgraphs of a larger dataset twitch-gamers. We report characteristics for all of them since they have different sizes, edge density, and may have different structural properties. The labels indicate if a streamer uses explicit language. Genius [19], twitch-gamers [36], arxiv-year [14], snap-patents [18], and wiki [20] are recently proposed large-scale heterophilous datasets. For the wiki dataset we remove all isolated nodes. Roman-empire, amazon-ratings, minesweeper, workers, and questions [33] are recently proposed mid-scale heterophilous datasets. We additionally construct one more binary classification graph — twitter-hate. This graph is based on data from Hateful Users on Twitter dataset on Kaggle. [^6] https://www.kaggle.com/datasets/manoelribeiro/hateful-users-on-twitter The labels indicate if a user posts hateful messages or not. We remove all the unlabeled nodes from the graph and use the largest connected component of the resulting graph.
@@ E.2 Dataset characteristics
As for label informativeness, it can be seen that on real heterophilous datasets it is typically very low (with the exception of roman-empire dataset). This is in contrast to synthetic datasets used for experiments in [22, 24], which sometimes exhibit a combination of low homophily and high label informativeness. High label informativeness of these datasets leads to strong GNN performance on them despite low homophily levels.
@@ F.1 Experimental setup
We use GCN [16], GraphSAGE [10], GAT [43] and Graph Transformer (GT) [40] as representative GNN architectures for our experiments. For GraphSAGE, we use the version with the mean aggregation function and do not use the node sampling technique used in the original paper. For all models, we add a two-layer MLP after every graph neighborhood aggregation layer and further augment all models with skip connections [11], layer normalization [2], and GELU activation functions [12]. For all models, we use two graph neighborhood aggregation layers and hidden dimension of 512. We use Adam [15] optimizer with a learning rate of $3\cdot 10^{-5}$ and train for 1000 steps, selecting the best step based on the validation set performance. We use a dropout probability of 0.2 during training. Our models are implemented using PyTorch [31] and DGL [44].
@@ F.2 Synthetic data based on SBM model
In Figure 2(a) of the main text, we show the results for GraphSAGE. In Figure 3, we additionally plot the dependence of GraphSAGE accuracy on homophily and label informativeness. It can be seen that the model’s accuracy is much more correlated with label informativeness than with homophily. The results for GCN, GAT, and GT are presented in Figures 4, 5, and 6, respectively. As can be seen, the performance of all models generally follows the same pattern, and $\mathrm{LI}$ is a better predictor of model performance than homophily. In particular, when $\mathrm{LI}$ is high, all models achieve high accuracy even if homophily is negative. Recall that we provide Spearman correlation coefficients between model accuracy and adjusted homophily or $\mathrm{LI}$ in Table 3.
F.3 Semi-synthetic data from [24]
We use the same experimental setting described above for training GCN, GraphSAGE, GAT, and GT on modifications of the cora and citeseer graphs from [24]. The results for GraphSAGE and cora are provided in Figure 2(b) of the main text. The results for GraphSAGE and citeseer are shown in Figure 7.
@@ F.4 Synthetic data based on LFR model
Figure 7: Accuracy of GraphSAGE on semi-synthetic citeseer graphs from [24]
F.5 Synthetic data from [22]
@@
Figure 8: Label informativeness depending on edge homophily on synthetic graphs from [22]
The authors investigate how different levels of homophily affect GNN performance. They find that the curve showing the dependence of GNN performance on edge homophily (as well as on node homophily) is $U$-shaped: GNNs show strong results not only when edge homophily is high, but also when edge homophily is very low. Our label informativeness explains this behavior. We use the same data generating process as in [22] and find that the curve of label informativeness depending on edge homophily is also $U$-shaped (see Figure 8). Thus, on the datasets from [22], GNNs perform well exactly when label informativeness is high regardless of edge homophily. The $U$-shape of the label informativeness curve is not surprising: when edge homophily is very low, knowing that a node has a neighbor of a certain class provides us with information that this node probably does not belong to this class.